22. You are aware of a survey that shows that 9% of the population in your city suffers from sickle cell disease, an autosomal recessive disorder. From this information deduce the following:
i. Value of q:
ii. Value of p:
iii. Proportion of Carriers (heterozygotes):
Answer #1

Given -

Frequency of homozygous recessive (aa) = 9%

means q2 = 0.09

So q = 0.3

We know that-

p+q = 1

So p = 1-0.3 = 0.7

Now

Heterozygotes = 2pq = 2 × 0.7 × 0.3 = 0.42